What are the concentrations of H+, HSO4, SO4 , H2SO4 in a 0.20 M solution of sulphuric acid?K2 = 10^-2 M

Grade:11

1 Answers

Arun
6 years ago
HSO4- H+ + SO42- 

Ka = [H+][SO42-]/ [HSO4-] = 1.26 x 10^-2 = x^2 / 0.20-x 
solving by quadratic formula 
0.00252 - 1.26 x 10^-2 x =x^2 

x^2 + 1.26 x 10^-2 x - 0.00252=0 

x = - 1.26 x 10^-2 + sq.rt (0.000159 + 0.0101)/2 =0.044 = [SO42-]= [H+] 

[HSO4-]= 0.20 - 0.044 =0.16 M